OpenStudy (anonymous):

What is the distance between points (6, 5) and (–12, 17)? Round to the nearest tenth.

OpenStudy (amistre64):

subtract the points and pythag the results

OpenStudy (amistre64):

(–12, 17) -(6, 5) ---------- -18, 12; sqrt(18^2 + 12^2)

OpenStudy (anonymous):

aprox. 21 i think

OpenStudy (anonymous):

sqrt{(6+12)^2+(5-17)^2)} = sqrt{18^2+(-12)^2}= sqrt(324+144)=21.6333076527

OpenStudy (anonymous):

8.9 14.7 18.2 21.6

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

21.6

OpenStudy (amistre64):

well its a good thing we came close

OpenStudy (anonymous):

okay thats what i was thinking too. THANK YOU SO MUCH! haha

OpenStudy (anonymous):

\[\sqrt{(-12-6)^{2}-(17-5)^{2}}=\sqrt{(-18)^{2}+(12)^{2}}=\sqrt{324+144}=\sqrt468{}=21.6\]